Imagine that you are working at a lab station (at the Mountain Dew factory) with an ultra pure water sample. Unfortunately, your lab neighbor accidentally splashes a few small drops of HCl in your sample when you aren't looking. What is the new pH of your ultra pure water if the HCl concentration is 5.0×10−7M solution of HCl?
Answer #1

ultra pure water is neutral.

thus it has a pH of 7

but, accidentally there is HCl added ie 5.0 x 10^-7

since, HCl is a strong acid.. so dissociate completely to produce H+ ions.

thus H+ ions = 5.0 × 10^-7

and pure H2O has = 10^-7

thus total [H+] = 5× 10^-7. +. 10^-7

= 6 x 10^-7

pH = -log[H+]

= -log(6× 10^-7)

= 6.22